OpenStudy (anonymous):

Solve. -8.1=d-2.7 Enter only the value of the variable.

OpenStudy (anonymous):

-8.1=d-2.7 -8.1+2.7=d-2.7+2.7 d=-5.4

OpenStudy (anonymous):

d=-8.1+2.7=-5.4

OpenStudy (anonymous):

-5.4